# The syzygy distinguisher

### v4

@@ 5 The distinguisher, with and without shortening
The shortened code ${\mathsf{C}}_{k-s^{*}-r}$ has dimension $s^{*}+r$ and length $n-k+s^{*}+r$. As $s^{*}$ and $r$ are $o(n)$ we have $s^{*}+r\ll n-k+s^{*}+r$, and $\operatorname{d}_{GV}$ is close to the length. Moreover $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}<r$ by hypothesis. Thus Heuristic 1 applies, and we expect $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})=0$ with high probabilty when ${\mathsf{C}}$ is random. On the other hand, (91) gives $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})>0$ for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$. Using Algorithm 1 we obtain $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $i\leq r$, where $\mathbf{M}_{i}$ has size $(s^{*}+r)\beta_{i-2,i-1}({\mathsf{C}}_{k-s^{*}-r})\times\binom{s^{*}+r+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{k-s^{*}-r})$. As the rate $\frac{s^{*}+r}{n-k+s^{*}+r}$ goes to zero, by Heuristic 1 again we can estimate the largest dimension of this matrix as $\binom{s^{*}+r+1}{2}\operatorname{ind}(\varphi_{i-2,{\mathsf{C}}_{k-s^{*}-r}})\leq(s^{*}+r)^{3}\binom{s^{*}+r}{r-3}$, for random codes — and we can assume this holds also for codes in $\mathcal{T}$, otherwise computing this dimension readily provides a distinguisher with lower complexity.
Now for a concrete choice of $s^{*}$ and $r$, set
  • $e\;=\;\lfloor\log_{q}(t-1)\rfloor\;=\;\log_{q}(n)-\log_{q}\log_{q}(n)+O(1)$
  • $f\;=\;(e+1)t-\frac{q^{e+1}-1}{q-1}\;=\;et+O(t)$
so
By Theorem 3.1, for all $s$ and all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ we have $r_{\max}({\mathsf{C}}_{s})\geq f-s$, so we can take $s^{*}=k-f\approx R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n$ as claimed. (In the binary Goppa case we could use Corollary 4 instead of Theorem 3.1, but this leads to the same estimate on $f$, hence on $s^{*}$.) Then for $\varepsilon>0$, setting $r=\left\lceil(1+\varepsilon)\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\right\rceil\ll s^{*}$, we have $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}\approx\frac{(s^{*})^{2}}{n-k}\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\ll r$, and then we can let $\varepsilon\to 0$.
Last, for the complexity estimate (93), using $r\ll s^{*}$ again and Stirling’s formula, we get $\log_{q}\left((s^{*}+r)^{3}\binom{s^{*}+r}{r-3}\right)\approx r\log_{q}\left(\frac{s^{*}}{r}\right)$, and we conclude.
We observe it is possible to slightly relax the condition $m=\lceil\log_{q}(n)\rceil$: complexity remains subexponential as long as $\frac{m}{\log_{q}(n)}\to 1$ fast enough.
@@ 6 Conclusion and open problems
We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms, more precisely, subexponential in the error-correcting capability of the code. However this is only an asymptotic result. For concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the security parameter by a non-negligible factor.
@@
To what extent can the implementation of the distinguisher be improved? For instance, if we choose a monomial order in Algorithm 1, and have a reduced basis for one syzygy space, then the generalized Macaulay matrix constructed from it will be:
  • somewhat sparse
  • already partially reduced.
This could be exploited to make the computation of (a reduced basis of) the next syzygy space faster. More generally, optimizations used in related domains, for instance in Gröbner basis algorithms, could probably be imported.
(Interestingly, the author was led to introduce the shortening technique in this distinguisher by serendipity, after witnessing a strange phenomenon while partially implementing the optimization alluded in Problem 1.)
@@
Our lower bounds on the Betti numbers, and in particular the $r_{\max}$, of dual alternant or Goppa codes, are not tight in general. Can they be improved? Better, can one give a complete, explicit description of the minimal resolution of these codes? (See the Supplementary material for elements in this direction.) How would this impact the complexity of the distinguisher?
We initiated the study of syzygies of codes (or equivalently, of finite sets of points in projective space) from a genuinely coding theoretical perspective.
@@
Pursue this study in a systematic way. This should cover basic operations on codes that have a standard geometric interpretation, such as puncturing (discarding one subset from a set of points), shortening (projection from one or several of the points), duality of codes (“Gale duality”), tensor product of codes (Segre embedding), and powers of codes (Veronese embedding), but also other operations such as the $(u|u+v)$, subfield subcode, trace code, or concatenated code constructions; and crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.
@@
Can syzygies serve in decoding algorithms?
@@
Can our syzygy distinguisher be turned into a key recovery attack, or more properly, can some of the ideas beneath be used to build a key recovery attack? For instance, one would like to combine our results with:
@@
$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&222&1943&1725&\mathbf{1120}&\mathbf{700}&\mathbf{240}&\mathbf{35}\\ 1&193&1344&\mathbf{525}&\mathbf{420}&\mathbf{175}&\mathbf{30}&-\\ 2&165&801&\mathbf{225}&\mathbf{120}&\mathbf{25}&-&-\\ 3&138&312&\mathbf{75}&\mathbf{20}&-&-&-\\ 4&112&\mathbf{40}&\mathbf{15}&-&-&-&-\\ 5&87&\mathbf{10}&-&-&-&-&-\\ 6&63&-&-&-&-&-&-\\ 7&40&-&-&-&-&-&-\\ 8&18&-&-&-&-&-&-\end{array}$
@@
These Figures illustrate the last Experimental fact. Boldface values match (105). This strongly suggests that the minimal resolution of the codes contains $m$ conjugate Eagon-Northcott complexes of length $r^{*}$, whose top degree components are in direct sum.
@@
Figure 13 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $[56,16]_{2}$-codes. For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling). The average value of $\operatorname{def}(\varphi_{r})$ among these samples is displayed, and also its $99\%$ distribution interval (which means at most $0.5\%$ fall below and at most $0.5\%$ above).
Here we have $\frac{k(k+1)}{n}\approx 4.86$. Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable), while $\operatorname{def}(\varphi_{r})=0$ with high probability as we move away from these intervals.

### v5

@@ 5 The distinguisher
We use Proposition 8 with $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, and $r^{*}=f$ as provided by Theorem 3.1. The analysis will be easier if rephrased in terms of $k_{r^{*}}=k-r^{*}$ and $r=r^{*}-s$. First, Lemma 11 gives
$k_{r^{*}}\;=\;k-f\;\sim\;R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n.$ (93)
Let us then turn to condition (88). Fix an $\varepsilon>0$, and set
$r=\left\lceil(1+\varepsilon)\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\right\rceil\ll k_{r^{*}}$ (94)
and $s=r^{*}-r$. Then $k_{s}=k_{r^{*}}+r$ and $n_{s}=n-k+k_{r^{*}}+r$, so for $n$ large enough we get
$\frac{k_{s}(k_{s}+1)}{n_{s}}=\frac{(k_{r^{*}}+r)(k_{r^{*}}+r+1)}{n-k+k_{r^{*}}+r}\approx\frac{k_{r^{*}}^{2}}{n-k}\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n<r=r^{*}-s.$ (95)
Thus, for any $\varepsilon>0$, condition (88) is satisfied with this value of $r$ as soon as $n$ is large enough, say larger than some $n(\varepsilon)$. But this means precisely that as $n\to\infty$, we can let $\varepsilon=\varepsilon(n)\to 0$ while still satisfying the condition. We then get
$r\;\sim\;\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n.$ (96)
Moreover the shortened code ${\mathsf{C}}_{s}$ has rate $\frac{k_{s}}{n_{s}}=\frac{k_{r^{*}}+r}{n-k+k_{r^{*}}+r}=o(1)$, so by Remark 2 both conditions in Heuristic 1 are satisfied. Thus Proposition 8 applies.
[…]
In [5, §3.4], the security level $2^{b}$ — or less formally, the number $b$ of targeted “bits of security” — of the Classic McEliece system, essentially defined as the complexity of the currently best attack (namely, by information set decoding), is shown asymptotically to be linear in $t\propto n/\log(n)$. Then in (92) we have
$\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$
which means that the complexity of our distinguisher is subexponential in $b$ (although, admittedly, only very slightly so).
@@ 6 Conclusion and open problems
We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms; more precisely, subexponential in the error-correcting capability of the code. However this is only an asymptotic result: for concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the best attacks by a non-negligible factor.
@@
Improve the implementation of this distinguisher.333addendum: actually, the improvements discussed below can already be found in works of Kreuzer-Lusteck and Albano-La Scala
For instance, if we choose a monomial order in Algorithm 1, and have a reduced basis for one syzygy space, then the generalized Macaulay matrix constructed from it will be (i) somewhat sparse and (ii) already partially reduced. This could be exploited to make the computation of (a reduced basis of) the next syzygy space faster.
Alternatively, instead of iteratively computing a whole range of the linear strand, one could directly compute a targeted Betti number, e.g. using Koszul cohomology (the author would like to thank W. Castryck for this suggestion). It is then likely that this Koszul cohomology computation could also benefit from similar optimizations (i) and (ii) exploiting sparsity and structure.
@@
Our lower bounds on the Betti numbers, and in particular the $r_{\max}$, of dual alternant or Goppa codes, are not tight in general. Can they be improved? Better, can one give a complete, explicit description of the minimal resolution of these codes? (See at the beginning of the Supplementary material for elements in this direction.)
We initiated the study of higher syzygies of codes (or equivalently, of finite sets of points in projective space) from a genuinely coding theoretical perspective.
@@
Pursue this study in a systematic way. This should cover basic operations on codes (some of which have a nice geometric interpretation, some of which less so), and also crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.
@@
Can higher syzygies serve in decoding algorithms?
@@
Betti numbers actually are invariants of the monomial equivalence class of a code (Lemma 3). Can they find other applications, for instance, to problems of code equivalence?
Problem 7
Can our distinguisher be turned into a key recovery attack, or more properly, can some of the ideas beneath be used to build a key recovery attack? For instance, one would like to combine our approach with:
@@ 6.0.1 Acknowledgements
$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&222&1943&1725&\mathbf{1120}&\mathbf{700}&\mathbf{240}&\mathbf{35}\\ 1&193&1344&\mathbf{525}&\mathbf{420}&\mathbf{175}&\mathbf{30}&-\\ 2&165&801&\mathbf{225}&\mathbf{120}&\mathbf{25}&-&-\\ 3&138&312&\mathbf{75}&\mathbf{20}&-&-&-\\ 4&112&\mathbf{40}&\mathbf{15}&-&-&-&-\\ 5&87&\mathbf{10}&-&-&-&-&-\\ 6&63&-&-&-&-&-&-\\ 7&40&-&-&-&-&-&-\\ \end{array}$
@@
Figures 8-8 illustrate the last Experimental fact. Boldface values match (107). This strongly suggests that the minimal resolution of the codes contains $m$ conjugate Eagon-Northcott complexes of length $r^{*}$, whose top degree components are in direct sum.
Computing the Betti diagram of some families of codes
Regularity $2$ helps in computing the full Betti diagram of codes, as one row can be deduced from the other. In order to illustrate this, let us first recall that a $[n,k]$-code ${\mathsf{C}}$ is MDS if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})=k+1$.
Lemma 12 ([15, Th. 1], reformulated)
Proposition 10
$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (108)
Proof
Proposition 11
  • $\beta_{0,0}=1$
  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $12\leq r\leq k-$ 1
Proof
Now set $n=2k-1$.
Figures 10 and 10 illustrate Propositions 10 and 11.
Figure 9: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
Figure 10: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
Exercise 1
@@
Figures 13-13 compare Betti diagrams for various types of codes of length $n=23$ and dimension $k=12$. Observe that these parameters give $\frac{k(k+1)}{n}\approx 6.78$.
First, Figure 13 presents the Betti diagram as predicted by the minimal resolution conjecture, i.e. with all defects equal to zero. It turns out that the minimal resolution conjecture is true for these parameters, so generic $[23,12]$-codes over an infinite field actually have this Betti diagram. Such codes are MDS, i.e. they have $d=12$ and $d^{\perp}=13$, so the intervals $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$ are empty, and Experimental fact 1 is trivially verified. Actually, codes with this Betti diagram can be found even over not-so-large finite fields, e.g. $q=5$ suffices (the author did not search for optimality). These codes are no longer MDS, but they still have $d,d^{\perp}\geq 7$ and again Experimental fact 1 is trivially verified. Moreover, the converse Experimental fact 2 is also verified as all defects are zero.
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&884&1397&1224&490&121&18&1&-&-\\ 2&-&-&4&44&234&820&1738&1888&1222&485&110&11\\ \end{array}$
Figure 12: a (pseudo)random $[23,12]_{2}$-code ($d=3$, $d^{\perp}=4$)
Figure 13: the $[23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
Second, Figure 13 presents the Betti diagram of the binary code whose generator matrix
$G=\left(\begin{array}[]{ccccccccccccccccccccccc}1&1&0&0&1&0&0&1&0&0&0&0&1&1&1&1&1&1&0&1&1&0&1\\ 0&1&0&1&0&0&0&1&0&0&0&1&0&0&0&0&1&0&1&1&0&1&0\\ &&&&&\vdots&&&&&&&&&&&&\vdots&&&&&\\ 0&1&0&0&1&0&1&0&0&0&1&0&1&0&0&1&0&1&0&0&0&0&0\end{array}\right)$
of size $12\times 23$, is constructed from the binary expansion of $\pi$. We consider this code as representative of random codes. One can check that it has $d=3$, $d^{\perp}=4$. Then Experimental fact 1 is verified, as the defect is positive for $4\leq r\leq 10$. Moreover at the extremities of this interval we have equality $\beta_{2,4}=A_{4}(C^{\perp})=4$ and $\beta_{9,10}=A_{3}(C)=1$, as predicted. And the defect is zero for $r<4$ and $r>10$ so Experimental fact 2 is verified also.
Last, Figure 13 reproduces Figure 3, the binary Golay code. As $d,d^{\perp}\geq 7$, Experimental fact 1 is trivially verified. On the other hand, Experimental fact 2 is not verified, but this only illustrates the fact that the Golay code is certainly not representative of random codes. Indeed it has a lot of algebraic structure, which explains it admitting special syzygies.
More systematically, Figure 14 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $[56,16]_{2}$-codes. For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling). The average value of $\operatorname{def}(\varphi_{r})$ among these samples is displayed, and also its $99\%$ distribution interval (which means at most $0.5\%$ fall below and at most $0.5\%$ above).
Here we have $\frac{k(k+1)}{n}\approx 4.86$. Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable); while $\operatorname{def}(\varphi_{r})=0$ with higher and higher probability as we move away from these intervals, in accordance with Experimental fact 2.
@@
Distinguishing Goppa codes with $q=4$, $m=4$, $t=4$, irreducible Goppa polynomial
Distinguishing Goppa codes with $q=2$, $m=6$, $t=3$, irreducible Goppa polynomial
What happens? It turns out the conditions in Heuristic 1 are not satisfied anymore, so we cannot ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$. Indeed, in order to have $\beta_{r-1,r}=0$, we need $r>k+1-d$, where $d=\operatorname{d}_{\min}({\mathsf{C}})$. For $r=7$ and $k=mt=18$ this gives $d\geq 13$. Then for smaller $d$, from Experimental fact 1, we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of ${\mathsf{C}}$. Recall the distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ with high enough probablity. Experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably. Now what happens is that, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
Remark 3
The sheer fact that the minimum distance is easily computable for codes with such small parameters actually provides a much more efficient alternative distinguisher. Indeed, turning back to the primal codes, the designed distance for binary Goppa codes with $t=3$ is $7$. Let us thus consider a distinguisher that decides that a given $[n,n-18]$-code is Goppa if it has $\operatorname{d}_{\min}\geq 7$, and that it is random otherwise. Experiments then show that, as long as $n\geq 32$, less than $2\%$ of random codes with these parameters have $\operatorname{d}_{\min}\geq 7$. This means that this distinguisher has more than $99\%$ success rate! And it remains remarkably effective even for smaller $n$. For instance, as low as $n=24$, less than half random codes have $\operatorname{d}_{\min}\geq 7$, so this distinguisher still has more than $75\%$ success rate.
In these experiments, random codes were defined by taking a uniformly random $(n-18)\times n$ generator matrix, which might quite often produce degenerate codes as $n$ approaches $18$. Other probability distributions preventing this problem could be used, which would slightly alter the success rate, but not the qualitative behaviour of this distance-based distinguisher.